Type yo Fama's Llamas has a weighted average cost of capital of 10.5 percent. The company's cost of equity is 17...

Type yo

Fama's Llamas has a weighted average cost of capital of 10.5 percent. The company's cost of equity is 17 percent, and its pretax cost of debt is 8.5 percent. The tax rate is 34 percent. What is the company's target debt-equity ratio? (Do not round your intermediate calculations.)


rev: 09_20_2012

1.2628
3.25
1.3957
1.3824
1.3292
Answer #1

.105 = w1*.17 + (1-.34)(1-w1)*.085

.105 = .17w1 + .66*.085 - .66*.085w1

(.17-.0561)w1 = 0.0489

w1 = .429324


debt to equity ratio = (1-w1)/w1 = 1.3292
Answer #2

equity to debt ratio is x

equity is x * debt

WACC = ( x * debt * 17 + debt * 8.5 * ( 1 - 0.34 ) ) / ( x * debt + debt )

10.5 * ( 1 + x ) = 17x + 8.5 * 0.66

10.5 + 10.5x = 17x + 5.61

6.5x = 4.89

x = 4.89 / 6.5


debt to equity is 1/x = 6.5 / 4.89 = 1.3292